# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v3

@@ Appendix A Proofs
To compare DDBI and DSBI, we first construct a unified error bound by analyzing the discrepancy between the interpolant-induced log density ratio and the true log density ratio. Let $(\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\sim\pi\in\Pi(q_{0}^{\prime},q_{1}^{\prime})$ for DDBI and $(\hat{\mathbf{X}}_{0}^{\prime},\hat{\mathbf{X}}_{1}^{\prime})\sim\pi_{2\gamma^{2}}$ for DSBI, where $\pi_{2\gamma^{2}}$ denotes the entropically OT coupling.
Let $q_{t}^{\prime}$ denote the interpolant-induced density at time $t$. The error functional $\mathcal{E}:=\left|\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\mathrm{d}t-\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|$ satisfies:
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]=\mathbb{E}_{\pi}\left[\left|\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\mathrm{d}t-\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|^{2}\right]\leq 2\underbrace{\mathbb{E}_{\pi}\left[\left|\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\mathrm{d}t\right|^{2}\right]}_{\mathcal{V}}+2\underbrace{\mathbb{E}_{\pi}\left[\left|\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|^{2}\right]}_{C_{0}},$ (42)
where $C_{0}$ is independent of the interpolant choice. Thus, error comparison reduces to analyzing the time-integrated variance $\mathcal{V}$. Expanding $\mathcal{V}$ via Fubini’s theorem:
$\mathcal{V}$ $=\mathbb{E}_{\pi}\left[\left|\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\mathrm{d}t\right|^{2}\right]=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{\pi}\left[\partial_{t}\log q_{t}^{\prime}\partial_{s}\log q_{s}^{\prime}\right]\mathrm{d}t\mathrm{d}s=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\partial_{s}\log q_{s}^{\prime}\big{|}\mathbf{X}_{s}^{\prime}\right]\right]\mathrm{d}t\mathrm{d}s$ (43)
$=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\big{|}\mathbf{X}_{s}^{\prime}\right]\right]\mathrm{d}t\mathrm{d}s\quad(\text{Markov property})$
$=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]+e^{-\frac{|t-s|}{\tau^{2}}}\left[\partial_{s}\log q_{s}^{\prime}-\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]\right]+\mathcal{O}(\tau^{4})\right]\right]\mathrm{d}t\mathrm{d}s\quad(\boldsymbol{\star})$
$=\int_{0}^{1}\int_{0}^{1}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]+e^{-\frac{|t-s|}{\tau^{2}}}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\right]\mathrm{d}t\mathrm{d}s+\mathcal{O}(\tau^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)\left(\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}[\partial_{s}\log q_{s}^{\prime}]\mathrm{d}s\right)+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\int_{0}^{1}e^{-\frac{|t-s|}{\tau^{2}}}\mathrm{d}s\right)\mathrm{d}t+\mathcal{O}(\tau^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\tau^{2}\left(2-e^{-\frac{t}{\tau^{2}}}-e^{-\frac{1-t}{\tau^{2}}}\right)\right)\mathrm{d}t+\mathcal{O}(\tau^{4})$
$\leq\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+2\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+\mathcal{O}(\tau^{4})\quad(\boldsymbol{\star}\boldsymbol{\star}).$
$(\boldsymbol{\star})$ The conditional expectation $\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\big{|}\mathbf{X}_{s}^{\prime}\right]$ satisfies the backward equation associated with the Markov semigroup. For small $\tau$, the process resembles an Ornstein-Uhlenbeck process near equilibrium, where the covariance decays exponentially with the time separation $|t-s|$ due to the Chapman-Kolmogorov equation.
$(\boldsymbol{\star}\boldsymbol{\star})$ This inequality holds because the term $2-e^{-\frac{t}{\tau^{2}}}-e^{-\frac{1-t}{\tau^{2}}}$ is strictly positive for $t\in[0,1]$ and attains its maxima at $t=0.5$ and minimum at $t=0$ and $t=1$. The first term is bounded by $C_{0}=\mathbb{E}_{\pi}\left[\left|\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|^{2}\right]$ according to the Fubini’s theorem and the inequality $\mathbb{E}[\mathbf{X}]^{2}\leq\mathbb{E}[\mathbf{X}^{2}]$.
Bringing this inequality back to Equation 42:
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]\leq 4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4}).$ (44)
To analyze $\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$, applying the law of total variance to $\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$:
$\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})=\mathbb{E}_{(\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\sim\pi}\left[\text{Var}\left(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right)\right]+\text{Var}_{(\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\sim\pi}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right).$ (45)
@@
The term $\mathbb{E}_{\pi}\left[\text{Var}\left(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right)\right]$. Taking the variance on both sides of Equation 46 and taking expectation over the coupling distribution $\pi$:
$\mathbb{E}_{\pi}\left[\text{Var}(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime}))\right]$ $=\mathbb{E}_{\pi}\left[\text{Var}\left(\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{Z}_{t}\right)+\text{Var}\left(-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}\right)+\text{Var}\left(-\frac{\dot{\sigma}_{t}^{2}\|\mathbf{Z}_{t}\|^{2}}{2\sigma_{t}^{2}}\right)\right]$ (47)
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}\right\|^{2}+0+\frac{\dot{\sigma}_{t}^{4}}{4\sigma_{t}^{4}}\text{Var}(\|\mathbf{Z}_{t}\|^{2})\right]$
@@
Specifically, when $\alpha_{t}=1-t$ and $\beta_{t}=t$, this reduces to $\mathbb{E}_{\pi}\left[\text{Var}(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime}))\right]=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}$.
@@
The term $\text{Var}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right)$. Taking the variance on both sides of Equation 46 and taking expectation over the coupling distribution $\pi$:
@@
$\text{Var}_{\pi}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right)$ $=\text{Var}_{\pi}\left(\mathbb{E}\left[-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}+\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{Z}_{t}+\frac{\dot{\sigma}_{t}^{2}\|\mathbf{Z}_{t}\|^{2}}{2\sigma_{t}^{2}}\right]\right)$ (48)
@@
$=\text{Var}_{\pi}\left(-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}+\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbb{E}[\mathbf{Z}_{t}]+\frac{\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}\mathbb{E}\left[\|\mathbf{Z}_{t}\|^{2}\right]\right)$
@@
$=\text{Var}_{\pi}\left(\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{0}+\frac{\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}d\right)=0.\quad\left(\mathbf{Z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{E}_{d}),\|\mathbf{Z}_{t}\|^{2}\sim\chi(d)\right).$
Bringing Equations 45, 47 and 48 into Equation 44, the expectational error bound with coupling $\pi$ becomes:
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ $\leq 2\mathcal{V}+2C_{0}=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4})$ (49)
@@
$=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\left[\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\right]\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4})$
$=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\int_{0}^{1}\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\mathrm{d}t\right]+4C_{0}+\mathcal{O}(\tau^{4}),$
@@
when $\alpha_{t}=1-t$ and $\beta_{t}=t$. Thus, the difference between the upper bound of $\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ for DSBI and DDBI becomes:
$\mathcal{V}_{\text{DDBI}}-\mathcal{V}_{\text{DSBI}}$ $=4\tau^{2}\underbrace{\left(1-e^{-\frac{1}{2\tau^{2}}}\right)}_{>0}\underbrace{\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]-\mathbb{E}_{\pi_{2\gamma^{2}}}\left[\left\|\hat{\mathbf{X}}_{0}^{\prime}-\hat{\mathbf{X}}_{1}^{\prime}\right\|^{2}\right]\right]}_{\geq 0(\boldsymbol{\star}\boldsymbol{\star}\boldsymbol{\star})}\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\mathcal{O}(\tau^{4}).$ (50)
$(\boldsymbol{\star}\boldsymbol{\star}\boldsymbol{\star})$ This inequality holds because $\pi_{2\gamma^{2}}$ is the solution to the entropically regularized OT problem:
$\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]-\mathbb{E}_{\pi_{2\gamma^{2}}}\left[\left\|\hat{\mathbf{X}}_{0}^{\prime}-\hat{\mathbf{X}}_{1}^{\prime}\right\|^{2}\right]$ (51)
@@
Finally, according to $\mathcal{V}_{\text{DDBI}}-\mathcal{V}_{\text{DSBI}}\geq 0$, the upper bound of $\mathbb{E}_{\pi_{2\gamma^{2}}}[\mathcal{E}_{\text{DSBI}}^{2}]$ is less than or equal to that of $\mathbb{E}_{\pi}[\mathcal{E}_{\text{DDBI}}^{2}]$.
@@
$\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]\leq\sqrt{\mathbb{E}_{q_{t}^{\prime}}[(\partial_{t}\log q_{t}^{\prime})^{2}]}=\sqrt{\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})+\left(\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\right)^{2}}.\quad(\text{Property of variance}).$ (52)
The first term on the r.h.s. is bounded according to Theorem 3.7, satisfying $\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}$ with $\sigma_{t}^{2}=t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon$. The second term on the r.h.s. vanishes identically $\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]=\int\frac{\partial_{t}q_{t}^{\prime}}{q_{t}^{\prime}}q_{t}^{\prime}\,\mathrm{d}\mathbf{x}=\int\partial_{t}q_{t}^{\prime}\,\mathrm{d}\mathbf{x}=\partial_{t}\left(\int q_{t}^{\prime}\,\mathrm{d}\mathbf{x}\right)=\partial_{t}(1)=0$. Thus, the term $\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]$ is bounded by:

### v5

@@ Appendix A Proofs
To analyze $\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$, applying the law of total variance to $\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$:
$\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})=\mathbb{E}_{(\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\sim\pi}\left[\mathrm{Var}\left(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right)\right]+\mathrm{Var}_{(\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\sim\pi}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right).$ (43)
@@
The term $\mathbb{E}_{\pi}\left[\mathrm{Var}\left(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right)\right]$. Taking the variance on both sides of Eq. 44 and taking expectation over the coupling distribution $\pi$:
$\mathbb{E}_{\pi}\left[\mathrm{Var}(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime}))\right]$ $=\mathbb{E}_{\pi}\left[\mathrm{Var}\left(\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{Z}_{t}\right)+\mathrm{Var}\left(-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}\right)+\mathrm{Var}\left(-\frac{\dot{\sigma}_{t}^{2}\|\mathbf{Z}_{t}\|^{2}}{2\sigma_{t}^{2}}\right)\right]$ (45)
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}\right\|^{2}+0+\frac{\dot{\sigma}_{t}^{4}}{4\sigma_{t}^{4}}\mathrm{Var}(\|\mathbf{Z}_{t}\|^{2})\right]$
@@
Specifically, when $\alpha_{t}=1-t$ and $\beta_{t}=t$, this reduces to $\mathbb{E}_{\pi}\left[\mathrm{Var}(\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime}))\right]=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}$.
@@
The term $\mathrm{Var}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right)$. Taking the variance on both sides of Eq. 44 and taking expectation over the coupling distribution $\pi$:
@@
$\mathrm{Var}_{\pi}\left(\mathbb{E}\left[\partial_{t}\log q_{t}^{\prime}(\mathbf{X}_{t}^{\prime}\mid\mathbf{X}_{0}^{\prime},\mathbf{X}_{1}^{\prime})\right]\right)$ $=\mathrm{Var}_{\pi}\left(\mathbb{E}\left[-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}+\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{Z}_{t}+\frac{\dot{\sigma}_{t}^{2}\|\mathbf{Z}_{t}\|^{2}}{2\sigma_{t}^{2}}\right]\right)$ (46)
@@
$=\mathrm{Var}_{\pi}\left(-\frac{d\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}+\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbb{E}[\mathbf{Z}_{t}]+\frac{\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}\mathbb{E}\left[\|\mathbf{Z}_{t}\|^{2}\right]\right)$
@@
$=\mathrm{Var}_{\pi}\left(\frac{\dot{\alpha}_{t}\mathbf{X}_{0}^{\prime}+\dot{\beta}_{t}\mathbf{X}_{1}^{\prime}}{\sigma_{t}}\cdot\mathbf{0}+\frac{\dot{\sigma}_{t}^{2}}{2\sigma_{t}^{2}}d\right)=0.$
The last equality holds because $\mathbf{Z}_{t}$ is a Gaussian noise, leading to $\mathbf{Z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{E}_{d})$ and $\|\mathbf{Z}_{t}\|^{2}\sim\chi(d)$.
Bringing Eqs. 45 and 46 into Eq. 43, the variance term $\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$ becomes:
@@
$\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})=\int_{0}^{1}\left[\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\right]\mathrm{d}t=\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\int_{0}^{1}\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\mathrm{d}t,$ (47)
@@
when $\alpha_{t}=1-t$ and $\beta_{t}=t$. Thus, the difference between the upper bound of the variance for DSBI and DDBI becomes:
$\mathrm{Var}^{\text{DDBI}}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})-\mathrm{Var}^{\text{DSBI}}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})$ (48)
$=$ $\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]-\mathbb{E}_{\pi_{2\gamma^{2}}}\left[\left\|\hat{\mathbf{X}}_{0}^{\prime}-\hat{\mathbf{X}}_{1}^{\prime}\right\|^{2}\right]$
@@
This inequality holds because $\pi_{2\gamma^{2}}$ is the solution to the entropically regularized OT problem. This completes the proof.
@@
$\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]\leq\sqrt{\mathbb{E}_{q_{t}^{\prime}}[(\partial_{t}\log q_{t}^{\prime})^{2}]}=\sqrt{\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})+\left(\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\right)^{2}}.\quad(\text{Property of variance}).$ (49)
The first term on the r.h.s. is bounded according to Theorem 3.7, satisfying $\mathrm{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}$ with $\sigma_{t}^{2}=t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon$. The second term on the r.h.s. vanishes identically $\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]=\int\frac{\partial_{t}q_{t}^{\prime}}{q_{t}^{\prime}}q_{t}^{\prime}\mathrm{d}\mathbf{x}=\int\partial_{t}q_{t}^{\prime}\mathrm{d}\mathbf{x}=\partial_{t}\left(\int q_{t}^{\prime}\mathrm{d}\mathbf{x}\right)=\partial_{t}(1)=0$. Thus, the term $\mathbb{E}_{q_{t}^{\prime}}[|\partial_{t}\log q_{t}^{\prime}|]$ is bounded by: